OpenStudy (anonymous):

You are walking on a beach and find a mystery piece of metal. You take it back to your lab and measure its mass to be 0.5 kg. In order to heat the metal 1 degree K, you must add 115 J of heat. Identify the metal. q =mCΔT

OpenStudy (zela101):

What does the variables stand for in q=m*C* delta t ?

OpenStudy (anonymous):

c=115/0.5=230 j/kg k but i think you should search for this metal that it's specific heat capacity is 230.